OpenStudy (anonymous):

(x-8) - 14 = -15 x =

OpenStudy (anonymous):

(x-8) - 14 = - 15 x-8-14 = -15 x - 22 = -15 x = -15 +22 x = 7

OpenStudy (anonymous):

thanks man @ashsek

OpenStudy (anonymous):

you are welcome @quentavess